Circular No. 6523

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1996bq, 1996br, 1996bs, 1996bz Christian Pollas, Observatoire de la Côte d’Azur (OCA), reports his

discovery of another apparent supernova in the field of SNe 1996bq, 1996br, and 1996bs (cf. IAUC 6502): SN 1996bz appears at α = 2^{h}31^{m}24^{s}.13, δ = +39^{o}43′51′′.3 (equinox 2000.0), which is 2′′.4 west and 4′′.6 north of the anonymous host galaxy. Estimated V magnitudes for SN 1996bz: Oct. 12 UT, 21.0: (faintly visible at limit of photograph); Nov. 3.9, ∼ 19.3; Dec. 15.87, 20.1: (faintly visible near limit of photograph). Nothing is present at the location of SN 1996bz on the original Palomar Sky Survey prints, on a recent Palomar Sky Survey B film, or on a 1995 OCA film (limiting mag ∼ 22.0); a nearby star of mag 20.0 is has position end figures 23^{s}.99, 48′′.4, and a second nearby star (V ∼ 15.5) has end figures 21^{s}.86, 53′′.9.

Pollas further notes that the Dec. 15.87 film also shows SN 1996bs at V ∼ 19.5 and SN 1996bq near the limit at V ∼ 20.0; SN 1996br is no longer visible.

SUPERNOVA 1996ca IN NGC 7300 S. Nakano, Sumoto, Japan, reports the discovery by Masakatsu Aoki,

Tsukioka Cho, Toyama, of a supernova (mag 16.5) on unfiltered CCD frames taken on Dec. 15.462 UT with a 0.43-m f/6 reflector. SN 1996ca is located at α = 22^{h}30^{m}59^{s}.26, δ = −13^{o}59′50′′.9 (equinox 2000.0), which is 8′′.7 west and 20′′.5 north of the center of NGC 7300. No star appears at this location on numerous patrol frames up through Dec. 3.

P. Garnavich and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that P. Berlind obtained a spectrogram of SN 1996ca on Dec. 17.1 UT with the 1.5-m Tillinghast telescope at the F. L. Whipple Observatory on Mount Hopkins; the spectrum shows this to be a type-Ia supernova, 3–5 days past maximum. Narrow emission lines from the host galaxy indicate a recession velocity of 5100 km/s.

SUPERNOVA 1996by IN UGC 3379 Weidong Li, Beijing Astronomical Observatory, reports the following

precise position for SN 1996by (cf. IAUC 6522): α = 5^{h}58^{m}24^{s}.96, δ = +68^{o}27′12′′.1 (equinox 2000.0).

1996 December 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green